Can I get a quik hand on these two,
if f(x)=5x-1 then how would i figure out f(a - 2)
If an inventor charges $4 per unit, then her profit for producing and selling x units is given by the function\
P(x)=2.25x-7000
a what is her profit after selling 2,000 units
b profit after 5,000 units
c. what would the break-even point for sales.
thank you ahead of time.

Answer by stanbon(75887)

if f(x)=5x-1 then how would i figure out f(a - 2)
f(a-2) = 5(a-2) -1
=5a-11
----------------
If an inventor charges $4 per unit, then her profit for producing and selling x units is given by the function\
P(x)=2.25x-7000
a what is her profit after selling 2,000 units
P(2000) = 2.25*2000-7000 = -2500 which is a loss of $2500
b profit after 5,000 units
P(5000) = 2.25*5000-7000 = $4259
c. what would the break-even point for sales.
Let P=0
Then 0 = 2.25x-7000
x= 3111.111...
Rounding down the break-even occurs when she produces 3111 units.
=============
Cheers,
Stan H.